by Angelika Drees
While rebucketing continues to look OK now we were fighting PS failure all night. We still want to get ready for delivering stores. Blue transmission deteriorated some though.
A series of problems beginning with blown up ramps due to BBQ kickers being on, a PS failure, fire alarms, exp. accesses and blown sweeps. This delayed rebuckting commissioning and store setup quite a bit. Therefore ramp/store development gets pushed into evening of Mar 24.
We had more ramp development and intensity increases during the past 2 days. First store with 48x48 over night. We had some problems with timing shifts in blue. Rebucketing not yet operational, longitudinal emittance needs to be addressed. This work is going on today. RF will test rebucketing, collimators will be moved in by the sequencer, gap cleaning will start automtically too. Pressure rise at IR4 gets close to critical with 48 bunches.
Lots of ramp and TF development over the past 3 days. Blue TF resulted in a reasonable rePlay ramp and good transmissions. Somehow yellow TF worked on a few ramps but didn't result in good transmissions. We are ramping yellow w/o rePlay. A lot of chromaticity changes on the ramp were made along with radius adjustments, orbit corrections and transition tuning. Intensity is up to a standard of 0.8 10^9 and today we had a ramp with 1.1 10^9/bunch an with 37x37 bunches. First collisions with 37x37 last night (Mar 19).
First optics measurements at store look good. During the day both bea transmissions got up to above 90%. Yellow TF ramp was successful and both ramps are rePlayed. Transmission efficiency deteriorated and we are working on that before increasing intensity. Ready for RF work at store (rebucketing).
BBQ work still going on. Blue transmission up to 95% and yellow around 50%. Since early morning we used rePlay to ramp blue while focussing on yellow. Exp. magnets were turned on today. Beam intensity rarely above .5 10^9.